Assume that a simple random sample has been selected from a normally distributed population and test the given claim. State the final conclusion that addresses the original claim and select three correct choices. A manufacturer makes ball bearings that are supposed to have a mean weight of 30 g. A retailer suspects that the mean weight is actually less than 30 g. The mean weight for a random sample of 16 ball bearings is 29.2 g with a standard deviation of 4.2 g. At the 0.05 significance level, test the claim that the sample comes from a population with a mean weight less than 30 g. Use the traditional method of testing hypotheses SELECT ALL APPLICABLE CHO A) There is NOT enough evidence to reject the Ho B) There is NOT enough evidence to support the Claim C) There is enough evidence to support the Ho D) Ho : μ<30 (claim) 1: H 30 E) Not enough evidence to reject the H1 F) Hi : μ<30 (claim) Ho 30 G) Not enough evidence to reject the Claim H) There is enough evidence to reject the Claim I) None of these
Answer #1

We know confidence interval is given by

\alpha =P[-z_\frac{\alpha}{2}<z<z_\frac{\alpha}{2}]

so,

0.05=P[-z_\frac{0.05}{2}<(\frac{\bar{x}-\mu}{\sigma})<z_\frac{0.05}{2}]\\ 0.05=P[-z_0.025<(\frac{29.2-\mu}{4.2})<z_0.025]\\ 0.05=P[-1.96*4.2<(29.2-\mu)<1.96*4.2]\\ 0.05=P[29.2-4.2*1.96<\mu<29.2+4.2*1.96]\\ 0.05=P[20.968<\mu<37.432]

from\ above \calculation \it\ is\ clear\ that\ rang\ of \mu\ is\ 20.968<\mu<37.432\\

Thus, there is not enough evidence to support or reject null hypothesis